write an equation in slope intercept form of the line that passes through (0,5) and (-3,-4)
Answer by palanisamy(496)

Let A be(0,5) and B be (-3,-4)
The equation of the line passing through the points A and B is
(x-0)/(0+3) = (y-5)/(5+4)
9x = 3(y-5)
9x = 3y-15
9x+15 = 3y
3y = 9x+15
y = (9/3)x+15/3
y = 3x+5